A shopkeeper has two types of coffee beans on hand. One type sells for $3.20 per lb, the other for $7.80 per lb. How many pounds of the $3.20 per lb coffee beans must be used in the mixture to produce 80 lbs of a blend that sells for $5.04 per lb?
Round your answer to 2 decimal places.



Answer by richwmiller(9143)

(3.2*x)+(7.8*y)=(80)*5.04
x+y=80
x=48 y=32